OpenStudy (anonymous):

find the fixed point f(x)=sqrt10+3x-4

OpenStudy (anonymous):

f(x)=2+4

jhonyy9 (jhonyy9):

inside sqrt is just10 ?

OpenStudy (anonymous):

no 10+3x is sqrt

OpenStudy (anonymous):

-4 is outside

jhonyy9 (jhonyy9):

so than sqrt(10+3x)=4 square both sides 10+3x=16 3x=6 x=2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

let me rewrite the function

jhonyy9 (jhonyy9):

go

OpenStudy (anonymous):

\[f\left( x \right)=\sqrt{10+3x}-4\]

OpenStudy (anonymous):

does this help

jhonyy9 (jhonyy9):

yes than make f(x)=0

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

and this is understandably ?

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

do I plug 0 in for x

jhonyy9 (jhonyy9):

not is zero check it what i have wrote before x=2

OpenStudy (anonymous):

oh ok

jhonyy9 (jhonyy9):

now is right all ?

OpenStudy (anonymous):

let me show you what I have

OpenStudy (anonymous):

\[10+3x=16\] 3x=6 x=2

OpenStudy (anonymous):

\[\sqrt{10+3(2)}-4\]\[\sqrt{10+6}-4\]\

OpenStudy (anonymous):

\[\sqrt{16}-4\]\[\sqrt{12}\]

OpenStudy (anonymous):

\[\approx3.46410\]

OpenStudy (anonymous):

am I totally off the wall?

jhonyy9 (jhonyy9):

not is right because you have said that 4 is outside of radical

OpenStudy (anonymous):

that is what I was thinking is what do I do with the 4

jhonyy9 (jhonyy9):

and sqrt16=4 because 4*4=16

OpenStudy (anonymous):

so back where I put sqrt 16-4 iswrong

OpenStudy (anonymous):

where is my mistake

jhonyy9 (jhonyy9):

yes there where you have wrote 10+3x=16 is right

OpenStudy (anonymous):

the rest is wrong?

jhonyy9 (jhonyy9):

from this resulted that 3x=6 so x=2

jhonyy9 (jhonyy9):

at finaly

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

than good luck bye

OpenStudy (anonymous):

bye thank you

jhonyy9 (jhonyy9):

yw

OpenStudy (zarkon):

if you want to find the fixed points you need to solve \[f(x)=x\]

OpenStudy (zarkon):

The answers are -3 and -2